OpenStudy (anonymous):

What is the solution to the given system? y = -6x -2 y + 2 = -6x

OpenStudy (anonymous):

@hartnn can you help? (:

hartnn (hartnn):

you have 'y' isolated in the first equation use this y and plug it in 2nd equation y = -6x-2 in 2nd equation gives you -6x -2 +2 = -6x can you simplify this further ?

OpenStudy (anonymous):

I combined the like terms (-6x) and got this -2 + 2 = 12x

hartnn (hartnn):

not actually, 12x will be +6x +6x whereas in your question, -6x -2 +2 = -6x we need to add 6x on both sides giving us -2+2 = -6x +6x now whats -6x +6x = ...?

OpenStudy (anonymous):

-12x?

OpenStudy (anonymous):

ooh wait blonde moment haha! it would be 0 right?

hartnn (hartnn):

nopes, when you factor out x from -6x+6x , we get x(-6+6) and whats -6+6 =... ?

OpenStudy (anonymous):

0

hartnn (hartnn):

yes! so, on right side we have a 0 on left side we have -2+2 which is ?

OpenStudy (anonymous):

0

OpenStudy (primeralph):

Same thing.

hartnn (hartnn):

right so we have 0=0 that is true FOR EVERY VALUE of x or y so, in such a case we say we have infinite solutions for x and y

OpenStudy (anonymous):

ohh I get it now!

hartnn (hartnn):

i'm glad to hear that :)